# Instance-level Performance Prediction for Long-form Generation Tasks

### v1

@@ Abstract
We motivate and share a new benchmark [^1] Benchmark URL forthcoming for instance-level performance prediction of long-form generation tasks having multi-faceted, fine-grained quality metrics. Our task-, model- and metric-agnostic formulation predicts continuous evaluation metric scores given only black-box model inputs and outputs. Beyond predicting point estimates of metric scores, the benchmark also requires inferring prediction intervals to quantify uncertainty around point estimates. Evaluation spans 11 long-form datasets/tasks with multiple LLMs, baselines, and metrics per task. We show that scores can be effectively predicted across long-form generation tasks using as few as 16 training examples Overall, we introduce a novel and useful task, a valuable benchmark to drive progress, and baselines ready for practical adoption today.
@@ 1 Introduction
Large language models (LLMs) can tackle increasingly challenging open-ended, long-form generation tasks across diverse modalities (e.g., text, images, code, etc.) Liu et al. (2024); Yin et al. (2024). Given the vast output space, many different generations of varying quality are possible Giulianelli et al. (2023). This motivates evaluation of long-form output quality that is: 1) fine-grained and nuanced (rather than binary correct/incorrect), and 2) multi-faceted, with diverse evaluation metrics used to assess different aspects of output quality Gao et al. (2025). For example, summarization quality might be measured via aspects such as brevity, fluency, coherence, consistency, and relevancy.
Like other AI models, LLMs are fallible, with varying output quality across different inputs Schellaert et al. (2025). Consequently, the ability to predict the quality of each output – across different inputs, tasks, models, and metrics – would have myriad benefits. Key among these is to guide appropriate trust and use for both human users Steyvers et al. (2025); Dhuliawala et al. (2023); Bhatt et al. (2021) and downstream agents Mielke et al. (2022); Yoffe et al. (2024). For example, outputs predicted to be of high quality might be used as-is, while those of lesser quality might be revised or discarded accordingly. Predicted quality could also provide a signal to optimize prompt selection or model parameterization Zhou et al. (2022), to select between different generations Chen et al. (2023), or to route inputs to whichever model is predicted to perform best Lee et al. (2024); Šakota et al. (2024); Lu et al. (2023); Shnitzer et al. (2023); Ding et al. .
Assuming evaluation metrics for fine-grained output quality are continuous $\in[0,1]$, we not only want to predict point estimates for metric scores, but also prediction intervals to quantify uncertainty around point estimates Shrestha and Solomatine (2006); Patel (1989); Geisser (1993). For example, given a point estimate prediction of 0.8, a tight 95% interval (say [0.79, 0.81]) would predict that output quality is quite close to the point estimate, while a loose 95% interval (say [0.6, 0.1]) would mean that quality may actually be much lower. Prediction intervals thus provide crucial uncertainty quantification for risk-averse settings by indicating probabilistic lower bounds on expected performance Harrington et al. (2019). We thus emphasize the importance of prediction intervals as well as point estimates for performance prediction.
In this work, we motivate and formulate the task of instance-level performance prediction for long-form generation tasks: predicting instance-level output quality – point estimates and prediction intervals – for continuous evaluation metrics. By design, our task formulation is task-, model-, and metric-agnostic: given an input-output pair for any black-box model, one must predict arbitrary evaluation metrics. The generality of this formulation supports broad use across the diverse, ever-changing landscape of LLM models and evaluation metrics for long-form output quality. We assume supervision is key: by observing examples of past input-output pairs and corresponding metric scores, one can learn a metric-specific prediction function. An ideal solution would achieve strong predictive performance with minimal supervision.
To promote research on instance-level performance prediction for long-form generation tasks, we create and share a new benchmark spanning twelve long-form tasks/datasets: text summarization, machine translation, code generation, and question answering. With different metrics often used to assess different aspects of output quality, our benchmark includes multiple evaluation metrics for most tasks. In regard to LLM coverage , we include outputs from Llama 3.2 ;other LLMs could be easily added later.
Our benchmark includes two baselines. One uses off-the-shelf confidence estimators (verbalized confidence and output consistency) to elicit weak features, then trains a classic regression model (e.g., linear regression, random forests) to predict the metric score from the features. The other baseline uses a reference-free LLM-as-a-Judge Li et al. (2024a) with in-context learning for textual regression Vacareanu et al. (2024) to directly predict the metric score from the input-output pair.
Our work demonstrates the feasibility of instance-level prediction of continuous evaluation metric scores and associated uncertainty intervals. We show that scores can be predicted with low RMSE and CRPS, across long-form generation tasks, and using as few as 16 training examples (Section 5.2). While the feature-based regression baseline described above is shown to perform well, the LLM judge baseline proves inconsistent. We also see that some confidence-based features are much more predictive than others. Taken together, our formulation, benchmark, and findings provide both a strong foundation for future research as well as baselines suitable for practical use today.
@@ 2 Related Work
Our formulation of performance prediction draws upon a large body of prior work across related areas: Confidence Estimation and Uncertainty Quantification (Section 2.1), Score and Benchmark Predictability (Section 2.2), Query Performance Prediction in information retrieval (Section 2.3), Quality & Metric Estimation in machine translation (Section 2.4), Textual Regression (Section  2.5), and LLM-as-a-judge & Reward Models (Section 2.6).
2.1 Uncertainty and Confidence Estimation
Our work is particularly motivated by confidence estimation (CE) Xiong et al. (2024) and uncertainty quantification (UQ) Vashurin et al. (2025); Xiao et al. (2022), which similarly predict how much each system output can be trusted. While terminology and definitions vary across different works, UQ (Smith, 2013) canonically seeks to quantify certainty regarding a given process, observation, or system output, while CE seeks to predict the probability that a system output is correct or not (Guo et al., 2017; Steyvers et al., 2025).
However, as Huang et al. (2024) write, “the correctness of long-form generation is not always either true or false but can be partially correct. Therefore, a single confidence score for long-form outputs is ambiguous: it can either imply ‘I am x% confident that the answer is 100% correct’ or ’I am 100% confident that the answer is x% correct.” Moreover, Ulmer et al. (2024) suggest estimation of LLM confidence intervals as future work.
Nevertheless, prior work has typically modeled long-form generation confidence via binary correctness by: i) mapping generations to a most similar multiple choice answer (Tjuatja et al., 2024), ii) binarizing partial-credit scores via thresholding (Kuhn et al., 2023; Lin et al., 2024; Band et al., 2024; Pedapati et al., 2024), iii) testing whether or not the output contains an answer (string) (Rajpurkar et al., 2016; Yu et al., 2018), iv) focusing on tasks with short answers wherein evaluation can rely on exact matching (Pedapati et al., 2024; Xiao et al., 2022); or v) decomposing a long output into smaller nuggets checked individually for correctness (Zhang et al., 2025, 2024a; Min et al., 2023).
In recent work, Zhang et al. (2025) usefully infer partial-credit scores for correctness at various granularities (e.g. sentence or passage level) but still assume that output quality can be measured via a single dimension of correctness. Huang et al. (2024) also emphasize “…the multifaceted nature of evaluating long-form outputs, where factors like factuality, coherence, clarity, and comprehensiveness each play a role…” However, they assume evaluation dimensions are implicit and latent, leading to subjective disagreements between evaluators. In contrast, we assume distinct evaluation metrics assess different aspects of output quality, and predict each separately. Prior studies also tend to show only correlation between CE vs. evaluation metric scores, while we predict metric scores directly.
Our formulation pursues uncertainty quantification via prediction intervals that quantify uncertainty around point estimates of metric scores.
2.2 Score & Benchmark Predictability
Like CE and UQ above, score predictability work Hernández-Orallo et al. (2022); Pacchiardi et al. (2025); Schellaert et al. (2025) similarly makes instance-level predictions of binary output correctness. Hernández-Orallo et al. note continuous prediction is more complex and proceed with binary. Schellaert et al.  suggest future non-binary work while Pacchiardi et al. assume the binary case.
Three other differences are noteworthy. First, score predictability emphasizes anticipatory prediction of instance-level success before output generation. This avoids computing outputs but makes prediction more difficult (without need for our use cases). Our performance prediction task more closely resembles UQ and CE in which confidence is inferred for a given output. Second, model details are often exploited to improve prediction, vs. our model-agnostic, black-box formulation. Third, score predictability methods tend to leverage a large feature set (e.g. text embeddings); we require only (input, output, score) triples.
Other work on benchmark prediction infers continuous scores, but only for entire datasets in aggregate Ye et al. (2023); Zhang et al. (2024b). Instead, we predict the quality of individual outputs.
2.3 QPP in Information Retrieval
Our nomenclature of “performance prediction” is inspired by query performance prediction (QPP) work in information retrieval (Cronen-Townsend et al., 2002; He and Ounis, 2006; Arabzadeh et al., 2024). Given an input query, QPP seeks to predict the quality of output search results. Like our work, QPP is both instance-based and predicts arbitrary, continuous-valued metrics of output quality.
QPP work is split into pre-retrieval vs. post-retrieval methods. Like anticipatory score predictability (Section 2.2), pre-retrieval methods require predicting output quality (of search results) before executing the (input) query. In contrast, post-retrieval methods run the query and analyze output search results to better predict their quality, and this helps them to perform much better in practice. We also expect output analysis to help prediction.
2.4 Quality and Metric Estimation in MT
Quality estimation (QE) in machine translation (MT) predicts translation output quality at run-time Specia et al. (2009, 2010). Earlier work on CE for MT also assumes binary correctness Blatz et al. (2004) which QE then generalizes to continuous, partial-credit evaluation. The term CE is also still sometimes used to refer to the broader QE task Chelba et al. (2020). QE is canonically framed as a regression task to predict human quality ratings (i.e., “direct assessment”) of output quality.
Scarton et al. (2016) frame QE more broadly as predicting output quality across diverse NLP tasks, including text summarization and simplification as well as QE. Some QE work has also emphasized the value of inferring prediction intervals beyond simple point estimates Glushkova et al. (2021); Giovannotti (2023); Zerva and Martins (2024).
Zouhar et al. (2023) propose “metric estimation”: predicting evaluation metric scores rather than human quality ratings. However, their goal in doing so is to pre-train a QE model. We are agnostic as to the type of prediction target (i.e., metric score or human rating) used to evaluate long-form outputs. To illustrate this, we also include a “direct assessment” task in our evaluation (Section 5.1), complementing our other metric prediction results.
2.5 Textual Regression
Wang et al. (2022) note “almost no work in NLP on calibration in a regression setting” and consider three regression tasks: semantic similarity, sentiment analysis, and QE. They use BERT/SBERT and report point estimates and prediction intervals.
LLM-based work has pursue three primary approaches: i) extracting LLM features for classic regression Tang et al. (2025a), ii) leveraging in-context learning Vacareanu et al. (2024); and iii) fine-tuning for specific regression tasks Tang et al. (2025b). Vacareanu et al. (2024) consider three synthetic regression tasks and show effective results across several models via in-context learning, but without prediction intervals. We report a similar in-context LLM baseline in this work (Section 4.3). Most recently, Vedula et al. (2025) propose quantile regression for LLM-based price estimation, reporting prediction intervals as well as point estimates.
2.6 LLM-as-a-Judge & Reward Models
The rise of LLMs has also brought the advent of LLM-as-a-judge. A key distinction exists between reference-free (RF) methods, suitable for predicting output quality at run-time, vs. reference-based (RB) methods that require human reference outputs to compare AI outputs against Sheng et al. (2024); Li et al. (2024a). Tan et al. (2024) present an excellent discussion relating RF methods to reward models used in reinforcement learning from human feedback (RLHF), noting “these two fields are closely related and are gradually converging.” Our work is related to both fields, though we focus on run-time prediction of RB metrics.
We also note similarity between RF LLM judges and “verbalized” CE methods (Section 2.1) in which an LLM is prompted for its confidence that a given output is correct (Tian et al., 2023; Xiong et al., 2024). Both support run-time use of LLMs through prompting, though verbalized CE methods have assumed a particular type of assessment (i.e., confidence) while RF LLM judges have been conceived and pursued much more broadly.
While RF LLM judges are increasingly being used as final evaluators, their reliability remains contested: they can exhibit self-preference and positional biases Wang et al. (2024); Li et al. (2024b), mis-rank instances they cannot themselves solve without a human reference answer Krumdick et al. (2025), and underperform or distort correctness in domains like software engineering Zhou et al. (2025). More broadly, RF methods have inherent structural biases—being effectively one generation model evaluating another—which risks test-time gaming, favoring models similar to the judge, and even disadvantaging higher-quality (including human) outputs Deutsch et al. (2021). Reproducibility can also be limited by underlying changes in closed LLM judges or prompt sensitivity in use of open judges Bavaresco et al. (2025).
Meta-evaluation of RF methods often reports correlation with human judgments, but caution is warranted. For example, the breadth and diversity of correlation studies is limited Bavaresco et al. (2025). Krumdick et al. (2025) note correlation can also mask disagreement on difficult instances, limiting ability to evaluate more advanced models or reliably predict instance-level quality. Dorner et al. (2025) show that when a RF LLM judge has comparable accuracy to an evaluated model, “no debiasing method can decrease the required amount of ground truth labels by more than half… [showing] severe limitations of the [RF] LLM-as-a-judge paradigm…" Their empirical results show practical performance is even worse than this theoretical limit. Gera et al. (2025) similarly report, “Crucially, even very good instance-level capabilities do not guarantee accurate model ranking.”
While we expect RF LLM judges to both continue to improve and become better understood, the limitations above suggest that RB LLM judges (and perhaps classic RB metrics) still remain an important part of the NLP evaluation toolkit for evaluating long-form outputs. This motivates a continuing need for run-time prediction of RB evaluation metrics, as we pursue in this study.
3 Task Formulation
Prediction intervals represent a natural extension of uncertainty for point estimates, bounding the range within which the true performance score will fall with a specified probability (e.g. 95 %) Shrestha and Solomatine (2006); Patel (1989); Geisser (1993). Prediction intervals are increasingly being reported, e.g., in MT quality estimation (Section 2.4) and other textual regression tasks (Section 2.5).

### v2

# PIE: Performance Interval Estimation for Free-Form Generation Tasks

@@ Abstract
Confidence estimation infers a probability for whether each model output is correct or not. While predicting such binary correctness is sensible for tasks with exact answers, free-form generation tasks are often more nuanced, with output quality being both fine-grained and multi-faceted. We thus propose Performance Interval Estimation (PIE) to predict both: 1) point estimates for any arbitrary set of continuous-valued evaluation metrics; and 2) calibrated uncertainty intervals around these point estimates. We then compare two approaches: LLM-as-judge vs. classic regression with confidence estimation features. Evaluation over 11 datasets spans summarization, translation, code generation, function-calling, and question answering. Regression is seen to achieve both: i) lower error point estimates of metric scores; and ii) well-calibrated uncertainty intervals. To support reproduction and follow-on work, we share our data and code. [^1] Data and code will be shared upon paper acceptance.
@@ 1 Introduction
Large language models (LLMs) are fallible: the quality of their outputs varies substantially across inputs, tasks, models, and evaluation metrics Schellaert et al. (2025). As LLMs are increasingly applied to real-world tasks, the ability to predict the quality of each individual output is important to guide appropriate trust in model outputs for both human users and downstream agents Steyvers et al. (2025); Dhuliawala et al. (2023); Bhatt et al. (2021); Mielke et al. (2022); Yoffe et al. (2024).
Confidence estimation (CE) canonically estimates the probability that a model output is correct or not. Predicting such binary correctness is sensible for tasks with exact answers, but modern free-form generation tasks are often more nuanced (e.g., summarization, translation, code generation, etc.). With such tasks, output quality often varies both 1) continuously; and 2) along multiple aspects (aka dimensions or properties), depending on the task. Example aspects might include properties such as fluency, relevancy (Liu et al., 2023), factual consistency (Min et al., 2023), etc. Predicting only unidimensional binary correctness can thus fail to capture more nuanced differences and other important dimensions of output quality.
Why represent output quality as continuous? First and foremost, fine-grained evaluation metrics typically yield continuous scores; we want to be able to predict any arbitrary metric. Second, while subjective human quality ratings may be labeled ordinally, more objective quality measures are naturally continuous (e.g., execution time of generated code or time required to further revise model outputs). Framing the quality prediction task as regression allows us to be agnostic as to both: 1) whether the prediction target is ordinal or continuous; and 2) whether we are predicting an observed value (e.g., human rating or program execution time) or a derived evaluation metric.
This regression framing also has important implications for uncertainty quantification (UQ). When quality is binary, we can formulate uncertainty as the probability of an output being correct or not. However, when output quality is continuous, UQ must also shift to estimating confidence intervals around point estimates of output quality.
Such UQ is particularly consequential in high-stakes settings where decisions are made directly from model outputs, necessitating risk-aware judgments. For example, consider software development workflows where LLM-generated code is increasingly executed or deployed with limited human oversight (Cotroneo et al., 2025; Navneet and Chandra, 2025). High confidence is required that generated code will reliably execute, pass unit tests, and be efficient and robust (Tong and Zhang, 2024) [^2] Predicting generated code quality pre-execution avoids both potential security risks and any time/space requirements.. To enable risk-informed decision-making, a point estimate of output quality alone is insufficient; a confidence interval helps lower-bound quality to better assess the level of risk.
For example, imagine generated code has predicted functional correctness of 0.89 with a narrow 95% interval (e.g., $\pm 0.09$), meaning 95% confidence that output quality is $\geq$ 80%. Assuming point estimates and confidence intervals are well-calibrated (i.e., trustworthy), this may be considered sufficiently low risk to use the output without human review. Alternatively, if the same predicted correctness of 0.89 were instead accompanied by a wide interval (e.g., $\pm 0.28$), this would indicate substantial uncertainty and a more serious worst-case scenario: i.e., 95% confidence that output quality is only $\geq$ 0.61. In practice, such uncertainty can translate into failures that are costly to detect and resolve, motivating additional safeguards, such as requiring human review and revision. Table 1 illustrates such an example in more detail. Prediction intervals are thus essential for guiding risk-aware decision making (Hullman et al., 2025). Intervals must also be well-calibrated, since miscalibrated intervals can do harm in leading users astray.
In sum, this work identifies the need for: 1) supporting arbitrary, task-specific measures of output quality for nuanced, free-form generation tasks; 2) continuous, multidimensional point estimates of output quality; and 3) confidence intervals around point estimates. To address these needs, we propose Performance Interval Estimation (PIE). The goal of PIE is to enable calibrated trust in model outputs for such tasks via a combination of predicting: 1) point estimates of continuous, task-specific quality measures; and 2) calibrated confidence intervals around each point estimate.
By jointly modeling quality and uncertainty, PIE supports practical deployment decisions such as abstention, asking clarification questions, or fallback to alternative models in low-confidence cases (Aliannejadi et al., 2019; Wen et al., 2025; Chuang et al., 2025). High-quality outputs may be accepted directly, while low-quality ones can be revised, regenerated, or discarded. Predicted quality could further support prompt or model selection Zhou et al. (2022), candidate ranking Chen et al. (2023), and routing inputs to models expected to perform best Lee et al. (2024); Šakota et al. (2024); Lu et al. (2023); Shnitzer et al. (2023); Ding et al. (2024).
For reproducible benchmarking on PIE’s prediction objective, we identify a set of 11 datasets spanning summarization, machine translation, function-calling, code generation, and question answering. Since different evaluation metrics capture different aspects of output quality, we include multiple metrics per task. To benchmark over outputs from different generative models, we include outputs from Llama 3.2-11B and Gemini 2.0-flash-lite.
We then evaluate two approaches: reference-free LLM-as-judge vs. classic regression using confidence estimation features. For the former, we prompt an LLM to directly predict the target evaluation metric score and a corresponding prediction interval for each generated output. Our evaluation includes use of in-context learning with a small set of labeled examples. The classic regression model predicts metric scores and intervals from confidence estimation signals, including verbalized confidence, consistency-based measures, and other uncertainty methods, by training a regression model to map these features to continuous metric targets.
Results across the 11 datasets, 15 metrics, and 2 LLMs show that the regression-based approach performs best in estimating both point estimates and confidence intervals for metric scores. Because free-form gold references can be expensive to collect, sample-efficiency is also important. Results show the regression-based approach converges after only 16 training instances across all 11 datasets.
Overall, we conceptually integrate three theoretically related but methodologically distinct lines of prior work: confidence estimation, quality estimation, and LLM-as-a-judge (Section 2). In reviewing each line’s core assumptions, methodological choices, and limitations, we show none provides a task-agnostic, general solution for predicting uncertainty in free-form generation. We then introduce a unified task formulation, PIE, to predict continuous evaluation metrics together with calibrated uncertainty intervals (Section 3). Since PIE is agnostic regarding evaluation based on directly observed values or derived performance metrics, it is equally suited to predicting both reference-based quality metrics (e.g., based on human ratings) as well as more objective metrics like code run-time, memory use, or tool call alignment.
To enable reproducible benchmarking, we collate and share a multi-task, multi-metric benchmark dataset (Section 4). Experiments (Section 5) and findings (Section 6) shows that classic regression matches or exceeds contemporary LLM-as-judge methods, consistently achieving more accurate and better-calibrated intervals while also using as few as 16 labeled examples to do so (Section 6).
Input Scale the input field to the range [0, 1] and display it as a DataFrame.
LLM output scaler = MinMaxScaler( feature_range=(0, 1))
scaled_values = scaler.fit_transform([[x] for x in l])
df = pd.DataFrame(scaled_values, columns=[’Scaled Values’])
return df
Target text scaler = MinMaxScaler()
l_scaled = scaler.fit_transform( l.reshape(-1, 1))
df = pd.DataFrame(l_scaled, columns=[’Scaled Values’])
return df
Prediction $0.8918\pm 0.2824$ at 95% probability
Gold score $0.7500$ (functional correctness)
Table 1: Illustrative PIE example on BigCodeBench. The CE-Reg predictor estimates functional correctness and a 95% prediction interval. This interval covers the gold score despite an optimistic point estimate, mitigating overconfident assessment of code that could fail for array-shaped inputs.
Table 2: Comparison of output quality estimation paradigms, illustrating their respective strengths and limitations in supporting continuous, multi-dimensional evaluation and uncertainty for free-form generation.
Confidence Estimation Quality Estimation Reference-free LLM-as-a-Judge Performance Interval Estimation
Used mostly in Limited (QA) Limited (MT) Various tasks Various tasks
Continuous targets No Yes Yes Yes
Multiple dimensions No Limited (error types) Yes Yes
Explicit uncertainty Yes Limited Limited Yes
@@ 2 Related Work
In this section, we review and compare key related work in confidence estimation (Section 2.1), quality estimation (Section 2.2), and (reference-free) LLM-as-judge (Section 2.3). Each of these varies in how uncertainty and metric prediction are considered. To relate these approaches and their relative capabilities, we compare them via several criteria in particular regard to evaluating the quality of free-form generations, as summarized in Table 2.
2.1 Confidence Estimation (CE)
Confidence estimation canonically models the reliability of a system output as the probability of an output being correct or not, producing a single scalar confidence value tied to binary correctness (Blatz et al., 2004; Guo et al., 2017; Steyvers et al., 2025). This view naturally fits short, discrete tasks with exact answers, such as factoid questions where correctness can be treated as all-or-nothing.
Two CE approaches are particularly well-known for their simplicity and applicability to black-box models: verbalized CE Xiong et al. (2024); Tian et al. (2023) and consistency-based CE (Kuhn et al., 2023; Lin et al., 2024). The former prompts an LLM to self-report its confidence alongside the generated output in a single LLM invocation. Consistency-based methods instead infer confidence from the degree of agreement across multiple independent generations. Computational overhead thus scales linearly with the number of samples.
As discussed earlier, framing output quality wrt. binary correctness is less suited to free-form, nuanced generation tasks in which assessment of output quality is often fine-grained (i.e., continuous), multi-dimensional, and task-dependent.
2.2 Quality Estimation (QE)
Quality estimation in machine translation (MT) predicts output translation quality at runtime Specia et al. (2009, 2010). In comparison to PIE, it shares both important similarities and notable differences. For example, early work on CE for machine translation also assumed binary correctness Blatz et al. (2004), which QE then generalized to continuous, partial-credit evaluation Specia et al. (2009, 2010). PIE is similarly inspired but extends this to diverse tasks, arbitrary metrics, and a strong emphasis on uncertainty and calibrated prediction intervals.
Work on QE has been largely restricted to MT, though other potential application areas have been discussed. For example, Scarton et al. (2016) suggested generalizing QE to tasks such as correcting grammar Napoles et al. (2016), generation Dušek et al. (2019), and others Specia et al. (2018). In contrast, PIE is fundamentally conceived as task- and domain-agnostic. Canonically, QE also refers to predicting human quality ratings, i.e., “direct assessment” judgments of translation quality. That said, some work has considered other metrics such as post-editing time Specia et al. (2018). In contrast, PIE is fundamentally metric-agnostic. While both QE and PIE are concerned with output “quality”, PIE’s notion of quality is broader, encompassing arbitrary task-dependent metrics of quality, such as the functional correctness of code.
Another key difference between QE and PIE lies in their respective treatments of uncertainty quantification (UQ). While recent work has explored uncertainty within QE, principally via conformal prediction (Glushkova et al., 2021; Giovannotti, 2023; Zerva and Martins, 2024), even here, uncertainty is not treated as a core modeling objective, and the scope remains confined to translation. In contrast, UQ is central and fundamental to PIE to support risk-aware decision making across diverse tasks and settings. As Table 1 illustrates, even a high predicted quality score can still present significant risk if the prediction interval is wide.
Within QE, Zouhar et al. (2023) proposed metric estimation: predicting reference-based evaluation metrics without access to the human reference. Similar to PIE, the focus is on predicting evaluation metrics, not just human judgments. However, like most QE work, its scope is translation-specific, focused on a small set of reference-based metrics, and predicts only point estimates.
2.3 LLM-as-a-Judge (LLMaaJ)
The rise of LLMs has also brought the LLM-as-a-judge paradigm. A key distinction is whether an LLMaaJ is used at runtime (online) to predict output quality vs. being used during evaluation (offline) to evaluate output quality. Closely related to this is the distinction between reference-based (RB) vs. reference-free (RF) methods. RB methods require a human (gold) reference output with which to compare the generated output and so are only suitable for use during evaluation (offline) (Sheng et al., 2024; Li et al., 2024). In contrast, RF methods have no such requirement, assessing output quality without comparing to any reference output (Liu et al., 2023). As such, RF methods are more flexible and can be used both online and offline.
We view RF-LLMaaJ as closely related to CE and QE since all three seek to predict output quality at runtime. However, a key limitation of LLM approaches seen in prior studies is their limited ability to regress continuous values, often struggling to produce calibrated intervals (Vedula et al., 2025), also seen in our own findings (Section 6).
RF-LLMaaJ shares conceptual similarities with “verbalized” CE methods (Section 4.1) that prompt a model for its confidence in its output correctness (Tian et al., 2023; Xiong et al., 2024). Both rely on prompting an LLM at runtime, but verbalized CE assumes a specific assessment target (confidence), whereas RF-LLMaaJ methods have been explored more broadly in evaluating model outputs.
RF-LLMaaJ methods also share conceptual similarities with QE methods (Section 2.2). As discussed, QE canonically seeks to predict human ratings (for translation quality) and only rarely to predict objective measures (e.g., post-editing time Specia et al. (2018)) or evaluation metrics Zouhar et al. (2023)). Similarly, RF-LLMaaJ work also typically seeks to predict human judgments and is evaluated by how well it aligns with them, rather than predicting reference-based evaluation metrics. As noted earlier, prediction of confidence intervals is atypical in QE work (Glushkova et al., 2021; Giovannotti, 2023; Zerva and Martins, 2024). Similarly, uncertainty quantification for continuous predictions (e.g., inferring confidence intervals) is also rare with LLMs (Hashemi et al., 2024).
As in CE and QE, a key premise of our work is that (offline) evaluation requires information not available (online) at runtime. Examples include human reference outputs or knowledge such as memory use by generated code once executed. The key challenge is thus online prediction of offline evaluation. However, when online RF-LLMaaJ is also the standard for offline evaluation, there is no need to predict it. We posit that despite growing popularity of evaluation via RF-LLMaaJ methods, many important evaluation settings will continue to require information not available at runtime.
3 PIE Task Formulation
As summarized in Table 2, existing methods either model uncertainty over binary correctness or predict continuous quality without task-general uncertainty estimates. We therefore propose PIE, which given a model output, predicts a continuous evaluation metric for it, along with a calibrated uncertainty interval. PIE extends point-estimate prediction by bounding the range within which the true performance score is expected to fall with a specified probability (e.g., 95%) Shrestha and Solomatine (2006); Patel (1989); Geisser (1993).